OpenStudy (anonymous):

If the planet Mars were orbiting the Sun in an orbit 3 times as big as current orbit, how many times longer would it take Mars to go around the Sun than it does now?

OpenStudy (anonymous):

What do you think? .. what does your gut or stomach say? ;-)

OpenStudy (anonymous):

Please elaborate

OpenStudy (anonymous):

I meant.. what does your intuition say?? without any math, what do you feel... could be the answer?

OpenStudy (anonymous):

this is exactly what i saw on the paper....... i felt the same way

OpenStudy (anonymous):

no i mean't.. what do you THINK IS THE ANSWER??!

OpenStudy (anonymous):

if i knew the answer, i wouldn't have asked

OpenStudy (anonymous):

I am asking can you make a guess.. by using some logic, then we can evaluate the answer and compare to what you guessed.. so u would know whether your logic was right.. lets begin with this question what do you think happens to the time taken .. do you think it increases? or decreases ? compared to before?

OpenStudy (anonymous):

the time taken for Mars to orbit around the Sun will increase?

OpenStudy (anonymous):

ok .. good.. now what do you think How much time does it increase? three times?? or more or less? after that.. i wanna ask you if you know any formula connecting Time of orbit and the Radius of orbit? any law ?? anything?

OpenStudy (anonymous):

Kepler's 1st law of the orbit? distance of foci/ 2x semi major axis time for orbiting the sun will be more than three times

OpenStudy (anonymous):

but the first law, doesn't say anything about orbit time right? which law connects Time of orbit (TIME PERIOD) and the orbital radius (SEMI MAJOR AXIS)

OpenStudy (anonymous):

law of the period?

OpenStudy (anonymous):

Maybe? what does the law of periods say?

OpenStudy (anonymous):

P^2/a^3

OpenStudy (anonymous):

P^2/a^3 = what?..

OpenStudy (anonymous):

time taken for a planet to orbit around the sun ^2 / semi major axis ^3

OpenStudy (anonymous):

well the law says SOMETHING ABOUT THAT RATIO right?? isn't your sentence incomplete?

OpenStudy (anonymous):

do you mean the square of the Period P is proportional to the cube of the semi major axis of its orbit? or do you mean the Newton's version of this law?

OpenStudy (anonymous):

yea yea.. what you said in the first sentence.. P^2 is proportional to a^3 hence P^2/a^3 = CONSTANT.. that value will remain THE SAMe.. which means if you change P, a will change Accordingly to keep that ratio same so now can you answer your question.. in your question, the mars has changed its orbital randius to three times.. so what should be new P.. such that the ratio remains the same?

OpenStudy (anonymous):

is it like p(old)^2/ a(old)^3 = p(new)^2/ a(new)^3?

OpenStudy (anonymous):

yes yes.. exactly like that.. now you get it..i am a sorry my net went off suddenly :-/

OpenStudy (anonymous):

sure no problem, thanks!